The table below contains the results of 400 employed adults where people were asked if they felt tense or stressed out at work

Gender Yes No
Male 89 111
Female 93 107

a. What is the probability that a randomly selected person will be male?

b. Given that the employed person felt tense or stressed out at work, what is the probability that the employed person was female?

c. Given that the employed person is female, what is the probability that she felt tense or stressed out at work?

d. Is feeling tense or stressed out at work independent of gender? Briefly explain.
Answer #1

Total number of employees = 400

Number of male = 89 + 111 = 200

a) The probability that a randomly selected person will be male

200 400

0,5

b) Number of people who felt tense or stressed out at work = 89 + 93 = 182

Given that the employed person felt tense or stressed out at work, the probability that the employed person was female

93 182

= 0.511

c) Number of female = 93 +107 = 200

Given that the employed person is female, the probability that she felt tense or stressed out at work

93 200

0.465

d) P( feeling tense or stressed out at work | female) = 0.511

and P( feeling tense or stressed out at work) 400 = 0.455

Here,

P( feeling tense or stressed out at work | female) eq P( feeling tense or stressed out at work)

So, we can write, feeling tense or stressed out at work is not independent of gender.